40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8225     Entry time: Mon Mar 4 19:52:03 2013
Author: Jamie, Yuta 
Type: Update 
Category: General 
Subject: input pointing mirror (TT1/TT2) control improved 

We improved the active tip-tilt (TT) controllers such that they now have filter banks at the PIT/YAW inputs, and at the coil outputs:

TT.png

This allowed us to do a couple of things:

  • normalize the matrix to unity, and move overall gains into the filters (we moved x100 gain into PIT/YAW)
  • slider control is now PIT/YAW OFFSET
  • potentially do coil balancing
  • allow for input excitiations
  • automatically record EPICS values

These are all big improvements.  The TT MEDM screens were appropriately updated.

We had to rebuild/restart c1ass, which reset the TT pointing.  We recorded all the values before hand and were able to recover the pointing easily.  Interestingly, there did appear to be hysteresis in pitch, which is maybe not entirely unexpected, but still worth nothing.

 

ELOG V3.1.3-